The starting five for a hockey team consists of two guards, two forwards, and one center. Two starting players will be seleNEED HELP ALL THIS QUESTIONS, PLEASE WRITE PROCESS, THANK YOU!
Answer #1

Answer:

Given,

Sample n = 2

A = 2 , B = 2 , C = 1

nCr = n! / (n-r)!*r!

a)

P(Two captains are both guards) = P(X = 2)

= 2C2 / 5C2

= 1 / 10

= 0.1

b)

P(At least 1) = P(X >= 1)

= [(2C1*2C1*1C0) + (2C1*2C0*1C1)] / 5C2

= [2*2*1 + 2*1*1] / 10

= 6/10

= 0.6

c)

P(X = 1) = 1C1 / 5C2

= 1/10

= 0.1

d)

P(Two captains play same position) = [2C2*2C0*1C0 + 2C0*2C2*1C0] / 5C2

= 2/10

= 1/5

= 0.2